### v2

# Exploiting Diffusion Prior for Generalizable Dense Prediction

Abstract
Contents generated by recent advanced Text-to-Image (T2I) diffusion models are sometimes too imaginative for existing off-the-shelf dense predictors to estimate due to the immitigable domain gap. We introduce DMP, a pipeline utilizing pre-trained T2I models as a prior for dense prediction tasks. To address the misalignment between deterministic prediction tasks and stochastic T2I models, we reformulate the diffusion process through a sequence of interpolations, establishing a deterministic mapping between input RGB images and output prediction distributions. To preserve generalizability, we use low-rank adaptation to fine-tune pre-trained models. Extensive experiments across five tasks, including 3D property estimation, semantic segmentation, and intrinsic image decomposition, showcase the efficacy of the proposed method. Despite limited-domain training data, the approach yields faithful estimations for arbitrary images, surpassing existing state-of-the-art algorithms. The code is available at https://github.com/shinying/dmp.
Figure 1: Generalized dense prediction. (left) We leverage the pre-trained text-to-image diffusion model  [57] as a prior for various dense prediction tasks. (right) With only a small amount of labeled training data in a limited domain (i.e., 10K bedroom images with labels) for each task, our method performs favorably against SOTA predictors [29, 7, 17] on arbitrary images.
1 Introduction
Text-to-image (T2I) diffusion models [57, 61, 21, 13] have achieved unprecedented progress on text-guided image generation, producing highly imaginative and realistic images from diverse and free-from textual descriptions. These advancements open up a new era of AI-aided content creation with applications spanning various domains [24, 49, 64, 72, 2, 59, 52]. However, the creativity of images generated by T2I models poses challenges for off-the-shelf dense (e.g., depth, normal, segmentation) prediction methods [7, 29, 17] due to the domain gap. For example, the ZoeDepth [7] approach fails to accurately predict the depth of the cubism painting shown in the sixth column in Figure 1. Such property predictions are vital for understanding high-level semantics of generated contents and can facilitate various downstream applications such as 3D imaging [65] and relighting [84].
Existing dense prediction models are typically trained on “real-world” images regardless of the training dataset scale. While these models aim for generalization, bridging the domain gap between real-world and T2I-generated images remains challenging, as we demonstrate in the right-hand side of Figure 1. A potential solution is to take advantage of the inherent generalizability of pre-trained T2I models, for example, by formulating dense prediction as image-to-image translation. However, while several recent efforts have been made to solve various image-to-image translations with pre-trained T2I models [49, 24, 9], we show in Section 4 that these methods are not directly applicable.
Leveraging pre-trained T2I models as a prior for dense prediction is challenging for two reasons. First, most dense prediction tasks are inherently deterministic, posing difficulties when adapting a pre-trained T2I model designed for stochastic text-to-image generation. Second, it is crucial to strike a balance between learning target tasks and retaining the inherent generalizability of pre-trained T2I models. In other words, learned dense predictors should generalize to arbitrary images from the training data in a limited domain.
In this paper, we propose DMP (Diffusion Models as Priors) to leverage the pre-trained T2I model [57] as a prior for generalized dense prediction. To resolve the determinism-stochasticity misalignment, we introduce a deterministic mapping between the input RGB images and output prediction distributions. Specifically, we reformulate the diffusion process as a chain of interpolations between input RGB images and their corresponding output signals, where the importance of input images gradually increases over the diffusion process. The reverse diffusion (i.e., known as denoising or generation in original T2I) process becomes a series of transformations that progressively synthesize desired output signals from input images. Without randomization, such as Gaussian noise imposed, the mapping is entirely deterministic. In addition, to retain the generalizability of the pre-trained T2I model while learning target tasks, we use low-rank adaptation [26] to fine-tune the pre-trained model with the aforementioned deterministic diffusion process for each dense prediction task. Figure 1 demonstrates the generalization ability of the proposed method on the deterministic normal, depth, and segmentation prediction problems.
We conduct extensive quantitative and qualitative experiments on five dense prediction tasks to evaluate the proposed DMP approach: 3D property estimation (depth, normal), semantic segmentation, and intrinsic image decomposition (albedo, shading). We show that with only a small amount of limited-domain training data (i.e., 10K bedroom training images with labels), the proposed method can provide faithful estimations of the in-domain and unseen images, especially those that the existing SOTA algorithms struggle to handle effectively. We summarize the contributions as follows:
  • We propose DMP, an approach leveraging the pre-trained T2I model as a prior for dense prediction tasks.
  • We design an image-to-prediction diffusion process that adapts the stochastic T2I model for deterministic dense prediction problems.
  • We use five dense prediction tasks to validate that the proposed method obtains faithful estimation on arbitrary images despite training with a small amount of data in a limited domain.
2 Related Work
Diffusion models. Diffusion models estimate a target data distribution by modeling the transition from a noisy version of the distribution [68, 25]. Recently, diffusion models have shown unprecedented quality in the text-to-image [57, 61] setting by training on large-scale datasets [63]. The advancements unleash various text-guided image manipulation applications [24, 49, 64, 2, 52, 11, 9, 46]. The stochasticity in the generation process, a preferred property in most applications, derives from the initial noise and additional noise added during the denoising process. However, the dense prediction problems are usually deterministic. Some recent methods adopt deterministic sampling algorithms [69, 33, 41] and reformulate the diffusion and generative process by $\alpha$-blending and de-blending [23]. Though the mapping between initial noise and outputs in target distributions is deterministic, the correlation between each noise and output sample is stochastic. If adapted to deterministic tasks, the model may generate high-quality outputs unfaithful to input images.
Image-to-image translation. The task aims to learn a mapping between two visual domains. Early efforts [90, 28, 35, 91, 48, 10] mostly make use of generative adversarial networks and cycle consistency loss to learn the translation from scratch. With StyleGAN [31, 32] showing high-quality synthesis on certain categories, some methods [74, 55] seek to achieve image-to-image translation using pre-trained StyleGAN by training an additional encoder. However, the translation is limited to certain categories. Recently, following the success of large-scale text-to-image models, attempts have been made to perform image-to-image translation with pre-trained diffusion models [24, 49, 9, 76]. These methods, however, are not directly applicable to the dense prediction task.
Fine-tuning text-to-image diffusion models. In addition to image-to-image translation, pre-trained diffusion models are adopted to take additional modalities [86, 39], be customized on certain objects [18, 58] or styles [87], and synthesize videos [42, 22, 19]. These methods train additional zero-initialized layers [86], manipulate attention modules [19, 39], learn a token embedding [18, 87], or learn parameter offsets with low-rank matrices [26]. In this work, we adopt the low-rank adaptation [26] to fine-tune only parameter offsets of attention layers.
Generative prior for dense prediction. Prior work has leveraged pre-trained generative models as a prior for other tasks, such as representation learning [15], as the latent features of pre-trained generative models are found to be rich in semantics [75, 83, 81]. Bhattad et al. [8] manipulates style latents of StyleGAN [30] and reveals its learned ability to estimate image properties, but the generalizability is limited. Some works reuse [3, 82, 88, 20] or merge [73] latent features of pre-trained denoising U-Nets to perform segmentation and depth estimation. Others [79, 78] transform generation models into multi-task generalists by standardizing the outputs of tasks as images. In this work, instead of developing a specific approach for a particular task, we focus on analyzing the potential of pre-trained models as a prior for general dense prediction through a universal transferring framework.
3 Method
Our goal is to leverage the pre-trained T2I diffusion model as a prior to learn a dense prediction task from a set of labeled training data $\mathcal{D}=\{(x^{i},y^{i})\}^{n}_{i=1}$, where $x^{i}$ denotes the input image, $y^{i}$ indicates the corresponding output (e.g., depth map), and $x^{i},y^{i}\in\mathbb{R}^{H\times W\times 3}$. We first describe the text-to-diffusion models used as the prior in Section 3.1. Then we introduce the proposed DMP approach in Section 3.2.
3.1 Text-to-Image Diffusion Models
We use the pre-trained T2I latent diffusion model [57] as the prior. It consists of an autoencoder and a U-Net. The autoencoder converts between an image $y\in\mathbb{R}^{H\times W\times 3}$ and its latent feature $\tilde{y}\in\mathbb{R}^{h\times\ w\times c}$, where $(h,w)=(H/8,W/8)$ and $c$ represents the channel size of latent features. Since we do not modify the autoencoder in the proposed approach, we use $y$ to represent the latent feature $\tilde{y}$ to simplify the annotation in this paper.
The U-Net model takes as input a text description and learns to reverse the following diffusion process that gradually turns the image $y$ into noise map $y_{T}$:
$y_{t}=\sqrt{\bar{\alpha}_{t}}y+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t}\hskip 14.22636pt\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (1)
where $t=[1,\cdots,T]$, and $\bar{\alpha}_{t}$ is the noise schedule [25].
3.2 Leveraging Diffusion Prior
There are two challenges to leverage (i.e., fine-tune) the pre-trained T2I approach for estimating the pixel-level output (e.g., depth) $y$ from an input image $x$: 1) determinism-stochasticity misalignment and 2) generalizability. We introduce the solutions to tackle these two issues as follows.
Figure 2: Deterministic diffusion process. We formulate the diffusion process as a chain of interpolations between an input image $x$ and output $y$. The U-Net model is fine-tuned to gradually transform the input $x$ to the desired dense prediction $y$.
Figure 3: 3D property estimation of arbitrary input images. The first row shows the input images, while the remaining rows present the normals and depth estimated by different approaches. The proposed DMP method gives faithful estimation, even on the images where the off-the-shelf [29, 7] schemes fail to handle.
Deterministic diffusion. The diffusion process described in Eq. (1) is designed specifically for stochastic image generation. However, the mapping between input images and outputs in dense prediction problems is typically deterministic. We observe that directly applying the diffusion process in Eq. (1) to the dense prediction tasks introduces unnecessary variation in outputs that leads to apparent artifacts. Therefore, we use the blending strategy [23, 40] and re-design the diffusion process as follows. Instead of converting from noise maps to images in the conventional T2I method, the diffusion process in our DMP approach directly maps between the input image $x$ and output $y$. As illustrated in Figure 2, the proposed diffusion process is formulated as
$y_{t}=\sqrt{\bar{\alpha}_{t}}y+\sqrt{1-\bar{\alpha}_{t}}x\hskip 14.22636ptt=[1,\cdots,T].$ (2)
As we can see from Figure 2 and Eq. (2), the proposed scheme gradually increases the weight of the input image $x$ over the diffusion process. This can be considered as progressively morphing the output $y$ into the input image $x$ via interpolation. As a result, we can fine-tune the U-Net model to reverse the diffusion process that interactively “de-morphs” the input image $x$ and gives the final prediction $y$.
Parameterization. We explore various parameterizations for the U-Net model to make different predictions. As discussed in Section 4.5, we empirically find that the v-prediction [62] works well for the dense prediction tasks. Specifically, the U-Net model $v_{\theta}$ is fine-tuned using the following mean square loss:
$L_{\mathrm{DMP}}=\mathbb{E}_{(x,y),t}\big{[}\|(\sqrt{\bar{\alpha}_{t}}x-\sqrt{1-\bar{\alpha}_{t}}y)-v_{\theta}(y_{t},t)\|^{2}_{2}\big{]},$ (3)
where $v_{\theta}(y_{t},t)$ is the U-Net prediction. The reverse diffusion process can then be formulated as
$y_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}(\sqrt{\bar{\alpha}_{t}}y_{t}-\sqrt{1-\bar{\alpha}_{t}}v_{\theta}(y_{t},t))$ (4)
$+\sqrt{1-\bar{\alpha}_{t-1}}x\hskip 19.91692ptt=[T,\cdots,1],$
where $y_{T}=x$ and $y_{0}$ is the desired output.
U-Net fine-tuning. To learn the target tasks while retaining the inherent generalization ability of the pre-trained T2I model, we use the low-rank approximation [26] scheme to fine-tune all the attention layers in the U-Net model to minimize the objective described in Eq. (3).
4 Experimental Results
4.1 Experiment Setup
We evaluate the proposed DMP approach using five dense prediction tasks, including 3D property estimation (i.e., depth and normals), semantic segmentation, and intrinsic image decomposition (i.e., albedo and shading).
Table 1: Quantitative comparisons on 3D property estimation. We compute the metrics using the estimated results and the pseudo ground truth generated by the off-the-shelf predictors.
Normal Depth
In-domain Out-of-domain In-domain Out-of-domain
[…]
SPADE [47] 0.0708 0.1635 0.1268 0.2833 0.2132 0.4961 0.1379 0.3587 0.3190 0.2554
DRIT++ [36] 0.0784 0.1723 0.1350 0.3006 0.3792 0.2458 0.2134 0.4374 0.2585 0.3212
SDEdit [44] 0.2599 0.5087 0.2675 0.5293 0.4656 0.3533 0.3240 0.6640 0.2495 0.3382
DDIB [71] 0.1849 0.4210 0.2271 0.4847 0.3087 0.5130 0.2367 0.6275 0.2788 0.3120
IP2P (hard) [9] 0.3017 0.5468 0.3168 0.5757 0.4834 0.3235 0.3358 0.6450 0.2252 0.3461
IP2P (learned) [9] 0.3550 0.7181 0.3397 0.6836 0.3965 0.3302 0.3494 0.5182 0.2664 0.3261
VISII [46] 0.2081 0.4386 0.2448 0.4895 0.3498 0.4405 0.2912 0.5364 0.2855 0.3181
DMP 0.0514 0.1156 0.0872 0.1886 0.1072 0.8861 0.1020 0.2117 0.6395 0.1360
Figure 4: Qualitative results. The first row shows the input images. In the following, every two rows show the results predicted by the off-the-shelf predictors (which we considered as pseudo ground truth) and those by the proposed method.
Datasets. To better analyze generalizability by varying training and test data domains, we conduct the evaluation with synthetic images. We first generate diverse text descriptions using a large language model [45] by filling a keyword in a prompt template modified from the one used in Parmar et al. [50]. We then use a text-to-image diffusion model [57] to synthesize the images according to the text descriptions. Second, we use the following off-the-shelf predictors to generate the pseudo ground truth for each image: Omnidata v2 [29] for surface normals, ZoeDepth [7] for monocular depth, EVA-2 [17] for semantic segmentation, and PIE-Net [14] for intrinsic image decomposition (albedo, shading). Finally, we follow the protocol used in Bhattad et al. [8] to generate a set of training data, and three sets of test data:
  • Training set: We generate $10$K labeled images using the keyword “bedroom”.
  • In-domain test set: We generate $2$K labeled images using the keyword “bedroom”.
  • Out-of-domain test set: We use the $409$ category names of the indoor scenes in the SUN dataset [80] as the keywords to generate $2$K labeled images. The set is considered to be out-of-domain compared to the training set. Nevertheless, the off-the-shelf models that provide the pseudo ground truth still work well since the images belong to normal indoor scenes.
  • Arbitrary test set: We use random textual descriptions to generate the images. Since the generated images are almost free-form, the off-the-shelf models cannot provide proper predictions. Therefore, we consider the off-the-shelf approaches as compared methods and present only the visual comparisons.
Note that we present the quantitative results only on the in-domain test set for semantic segmentation due to the semantic label set constraint. In addition, we show qualitative comparisons of bedroom images using diverse styles (e.g., cyberpunk, comic) to understand the generalization ability.
Table 2: Quantitative comparisons on semantic segmentation. We compute the metrics using the estimated results and the pseudo ground truth generated by the off-the-shelf predictors.
Bed Pillow Lamp Window Painting Mean
[…]
SPADE [47] 0.8677 0.6370 0.5861 0.3473 0.3659 0.2084 0.6925 0.5627 0.5249 0.3826 0.6074 0.4276
DRIT++ [36] 0.8485 0.4587 0.2427 0.1435 0.1218 0.0776 0.3023 0.2414 0.2579 0.2114 0.3546 0.2265
SDEdit [44] 0.0958 0.0901 0.3824 0.0864 0.1522 0.0651 0.4501 0.2593 0.1333 0.0746 0.2428 0.1151
DDIB [71] 0.3984 0.3040 0.2256 0.0637 0.1630 0.0593 0.4741 0.2896 0.1728 0.0881 0.2868 0.1609
IP2P (learned) [9] 0.0714 0.0620 0.0086 0.0042 0.0228 0.0116 0.3532 0.1699 0.0386 0.0192 0.0989 0.0534
VISII [46] 0.0060 0.0059 0.0261 0.0136 0.0014 0.0011 0.2576 0.1772 0.0013 0.0012 0.0585 0.0398
DMP 0.8947 0.8506 0.5871 0.3645 0.6399 0.4414 0.8338 0.7335 0.7490 0.6735 0.7409 0.6127
Figure 5: Qualitative results on semantic segmentation. The first, second, and third rows respectively show the input images, pseudo ground truth predicted by an off-the-shelf model, and our results. The out-of-domain samples in (b) are bedroom images in diverse styles.
Compared methods. We compare our method with the GAN-based image translation methods SPADE [47] and DRIT++ [43, 36]. These models are trained from scratch using the training set (i.e., $10$K labeled bedroom images). We also include for comparison the following approaches that leverage the pre-trained T2I model as the prior:
  • SDEdit [44]: We fine-tune using the training label images $\{y^{i}\}^{10\mathrm{K}}_{i=1}$ with the standard diffusion process in Eq. (1). Then we follow the original SDEdit approach that adds the noise to an input $x$ and uses the fine-tuned model to de-noise for generating the output $y$.
  • DDIB [71]: We use the same fine-tuned model in SDEdit and adopt the DDIB method to predict the output $y$ from the input $x$.
  • InstructPix2Pix (IP2P) [9]: We evaluate two versions. In IP2P (hard), we use the pre-defined instructions such as “make it into the corresponding depth map” as the input to the model for inference. For the second version IP2P (learned), we optimize the token $\ast$ in the input text “make it into $\ast$” using the training set.
  • VISII [46]: We use their approach for fine-tuning with the training set and inference.
4.2 3D Property Estimation
[…]
The quantitative comparisons are presented in Table 1, and qualitative results are shown in Figure 4. The proposed approach performs favorably against the compared algorithms in terms of accuracy and generalization ability. Although the $\mathrm{REL}$ and $\delta$ scores reported for our method degrade from the in-domain to out-of-domain test sets, the $\mathrm{RMSE}$ score of the relative depth remains the same. This is due to the scene scale difference between the training images (bedroom) and test images (larger spaces such as sports fields). The $\mathrm{RMSE}$ measurement is computed based on relative depth, which is more resistant to the scene scale change. Consistent with the quantitative results, we observe that the depth predicted by our method is in the correct order for the out-of-domain images in Figure 4.
Generalization ability. We demonstrate the generalizability of the proposed approach in Figure 3, where we use arbitrary images as the inputs. Although our method is fine-tuned with only $10$K bedroom images with labels, it faithfully estimates the 3D property even on those images that the off-the-shelf methods fail to handle.
Table 3: Quantitative comparisons on intrinsic image decomposition. We compute the metrics to measure the difference between the estimated results and the pseudo ground truth created by the off-the-shelf predictors.
Albedo Shading
In Out In Out
SPADE [47] 0.0021 0.0030 0.0031 0.0040
DRIT++ [36] 0.0296 0.0392 0.0309 0.0408
SDEdit [44] 0.0375 0.0471 0.0501 0.0671
DDIB [71] 0.0411 0.0443 0.0403 0.0557
IP2P (hard) [9] 0.0329 0.0479 0.0361 0.0421
IP2P (learned) [9] 0.0215 0.0250 0.0290 0.0309
VISII [46] 0.0145 0.0246 0.0275 0.0285
DMP 0.0041 0.0064 0.0051 0.0070
Table 4: Effect of different parameterizations. We fine-tune the U-Net model to predict different signals in each reverse diffusion step to get the final outputs: predicting the input image $x$, predicting the output $y$, and v-prediction described in Eq. (3). The experiment is conducted on the surface normal prediction task.
In-domain Out-of-domain
[…]
Predicting $x$ 0.0736 0.1629 0.1319 0.2764
Predicting $y$ 0.0590 0.1291 0.0888 0.1914
v-prediction 0.0514 0.1156 0.0872 0.1886
Figure 6: Qualitative comparisons between various parameterizations. We fine-tune the U-Net model to predict different signals in each reverse diffusion step to get the final output: predicting the input image $x$, predicting the output $y$, and v-prediction described in Eq. (3).
4.3 Semantic Segmentation
Semantic segmentation [89] is a fundamental visual understanding task. Since the prediction is categorical, we use a simple conversion for regression models to predict discrete labels. We first generate ground-truth labels using the off-the-shelf EVA-2 [17] model. The label maps are then transformed into color maps where each category has different colors. The training and inference of the diffusion model are conducted using the color maps (in the RGB space). During inference, the predicted color maps are converted to categorical label maps by assigning each pixel to its nearest category in the color space.
We report the intersection over union (IoU) and accuracy to measure the in-domain performance in Table 2. The prediction by the proposed scheme is reasonable across all categories compared to the existing methods. Figure 5 demonstrates the results of our method and pseudo ground truth with the in- and out-of-domain (i.e., bedroom images in diverse styles) input images. Particularly in out-of-domain examples, our model gives favorable predictions compared to the off-the-shelf approach, e.g., the painting of the first, the curtain of the second, the window of the third, and the carpet of the fourth images in Figure 5 (b). This validates our idea that leverages the pre-trained T2I diffusion model as the prior for better generalization.
4.4 Intrinsic Image Decomposition
Intrinsic image decomposition [6] recovers albedo and shading properties from RGB images. It facilitates applications such as recoloring [5] and relighting [66]. Similar to PIE-Net [14], we use the mean square error $\mathrm{MSE}$ as the evaluation metric to compute the distance between the predicted and pseudo ground-truth estimations. We showcase the qualitative results in Figure 4 and quantitative comparisons in Table 3. The reported errors of most methods are an order of magnitude larger than ours. Notably, while our method and SPADE work well in both in- and out-of-domain settings, we found our method less influenced by artifacts in pseudo ground truth and to offer more reasonable estimations. More details are provided in the supplementary document.
Figure 7: Effect of varying the numbers of diffusion steps. We report the in-domain performance on the surface normal prediction task. In the case of single-step, we train the U-Net model to directly predict outputs from input images.
Figure 8: Single vs. multiple diffusion steps. In the single-step approach, we train the U-Net network to directly predict outputs from input images. The single-step approach does not generalize to unseen images well.
4.5 Ablation Study
Parameterization. We study the effect of employing various parameterizations. Specifically, we fine-tune the U-Net model to make different predictions in each reverse diffusion step for obtaining the final prediction: 1) predicting the input image $x$, predicting the output $y$ (similar to $x_{0}$-prediction in standard diffusion models), and v-prediction described in Eq. (3). We formulate each parameterization in detail in the supplementary document. The quantitative results are shown in Table 4 and the qualitative comparisons are presented in Figure 6. Predicting the input image $x$ generates accurate results in the in-domain setting. However, it fails to generalize to unseen domains. On the other hand, predicting the output $y$ demonstrates preferred generalization capability, but tends to produce blurry results with few details. We choose to use the v-prediction approach as it produces accurate results of arbitrary images.
Number of diffusion steps. We analyze the performance of our method with different numbers of diffusion steps in the inference stage. The quantitative comparisons on the normal prediction are presented in Figure 7. Using $5$ steps strikes a good balance between inference speed and accuracy. Furthermore, we study an extreme case of using a single step. We train the U-net to directly predict the output $y$ from the input image $x$. Although it shows competitive performance in the in-domain setting according to Figure 7, we find that the performance degrades significantly with the arbitrary input images, as examples shown in Figure 8. Considering the generalizability, inference speed, and accuracy, we use $5$ generation steps for all the other experiments.
5 Conclusion
In this work, we leverage a pre-trained T2I diffusion model for generalizable dense prediction. The core of our method is the design of the deterministic diffusion process that adapts the stochastic T2I framework for the deterministic prediction tasks. With low-rank approximation, the proposed approach learns the target tasks while retaining the inherent generalization ability of the T2I model. We show that with only a small number of labeled training data in a limited domain (i.e., $10$K bedroom images), our DMP scheme makes faithful predictions on arbitrary images. We believe that this work establishes a foundation for achieving ultimately-generalizable visual understanding in the future.
Appendix A Parametrizations
As described in Section 3.2, we empirically find that parametrizing the U-Net model through estimating v-prediction [62] performs favorably against predicting inputs or outputs. We detail the formulation of predicting inputs and outputs as follows. The U-Net model $\hat{x}_{\theta}$ predicting inputs is fine-tuned with the mean square loss:
$L=\mathbb{E}_{(x,y),t}\big{[}\|x-\hat{x}_{\theta}(y_{t},t)\|^{2}_{2}\big{]},$ (5)
and the reverse diffusion process is formulated as
$y_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}\left(\frac{y_{t}-\sqrt{1-\bar{\alpha}_{t}}\hat{x}_{\theta}(y_{t},t)}{\sqrt{\bar{\alpha}_{t}}}\right)$ (6)
[…]
The U-Net model $\hat{y}_{\theta}$ predicting outputs is optimized with the loss function:
$L=\mathbb{E}_{(x,y),t}\big{[}\|y-\hat{y}_{\theta}(y_{t},t)\|^{2}_{2}\big{]},$ (7)
and the reverse diffusion process is
$y_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}\hat{y}_{\theta}(y_{t},t)+\sqrt{1-\bar{\alpha}_{t-1}}x$ (8)
$t$ $=[T,\cdots,1],$
Appendix B Additional Experimental Results
B.1 Reliability of Off-the-Shelf Estimators
We indicate that the off-the-shelf estimators are not always reliable, especially the approach for intrinsic image decomposition. We demonstrate with the example of albedo estimation in Figure 9 that the off-the-shelf approach generates apparent artifacts in shadow areas, such as corners or floors under beds. The approach fails to recover the correct albedo but instead generates black patches. Consequently, SPADE also learns this pattern, but our model tends to correct artifacts by performing accurate estimation, manifesting the ability of generalization.
Figure 9: Qualitative comparisons on albedo estimation. SPADE [48] and the off-the-shelf approaches generate artifacts in dark areas.
B.2 Real-World Evaluation
NYU Depth v2.
Following Ke et al. [34], we evaluate our method on NYU Depth v2 [67] according to the protocol of affine-invariant depth evaluation [54]. We generate prompts with BLIP-2 [37] to use the model trained on synthetic bedroom images. We scale and shift the depth predictions to align ground truths by solving least-square fitting. The comparison against other approaches is shown in Table 5. DMP performs comparably with some previous methods trained with large-scale data.
Table 5: Comparison of performance on NYU Depth v2 [67].
# Training Samples NYU v2
[…]
MiDaS [54] 2M 11.1 88.5
Omnidata [16] 11.9M 310K 7.4 94.5
DPT [53] 1.2M 188K 9.8 90.3
Painter [78] 24K 8.0 95.0
Marigold [34] 74K 5.5 96.4
DMP 10K 12.0 86.5
ADE20K.
We also investigate the performance of semantic segmentation with more classes, e.g. 150 classes in ADE20K [89]. We follow the encoding strategy proposed by Wang et al. [78] and convert class indices into 3-digit numbers with a $b$-base system, which can be represented in the RGB space. However, the performance is unsatisfying (lower than $20\%$ accuracy). With the number of classes increasing, the differences between colors are less distinguishable. The unlabeled areas, which can be simply ignored when calculating loss in the image space, become hard to tackle in the latent space. We leave the application of real-world many-class semantic segmentation for future exploration.
We conduct another analysis with a subset of ADE20K containing only images with beds. The train-test split is constructed by applying the same filtering to the original splits, resulting in 1825 training images and 189 test images. We also generate prompts with BLIP-2. The results are presented in Table 6. The performance across real and synthetic domains is similar, especially for large items.
Table 6: Comparison of performance between the subset of ADE20K [89] and the synthesized bedrooms.
Bed Pillow Lamp Window Painting Mean
[…]
B.3 Additional Ablation Study
Modeling.
We vary the trainable layers and the presence of text conditions when fine-tuning the model. Since the example tasks we choose in this work are not directly conditional on text, providing text descriptions might not be necessary. Accordingly, training cross-attention layers is optional. Table 7 shows that text condition and training cross-attention layers help improve the performance of in-domain samples, but the difference in out-of-domain samples is unnoticeable between the settings. This result suggests that we can adopt curated real ground truth datasets without text descriptions for training at the expense of a subtle performance drop. Alternatively, we can generate prompts with image captioning [37], which may lead to better performance.
Table 7: Analysis of training cross-attention layers and providing text condition. Both improve the performance of in-domain samples but make little difference in out-of-domain data.
In-domain Out-of-domain
[…]
Self-attn 0.0606 0.1290 0.0890 0.1871
Self-attn + text 0.0605 0.1293 0.0876 0.1844
All attn + text 0.0514 0.1156 0.0872 0.1886
Size of Training Data.
We analyze the effect of varying the size of training data. We compare fine-tuning the model for normal estimation with $100$, $1$K, $10$K, and $100$K generated bedroom images. As shown in Figure 10, increasing data size over $10$K improves little performance, so we conduct the other experiments with $10$K training images.
Figure 10: Quantitative performance of normal estimation with different sizes of training data.
Table 8: NYU Depth v2 [67] performance comparison of models trained with real and pseudo ground truths.
[…]
Hypersim [56] Real 13.0 85.0
Bedrooms Pseudo 12.0 86.5
Quality of Training Data.
We examine the influence of data quality by comparing the models trained with real and pseudo ground truth. We use Hypersim [56] as the real ground truth and evaluate the models with NYU Depth v2 [67]. Table 8 shows that there is no significant difference between the two models. The model trained with pseudo ground truth even performs slightly better. We speculate that the data diversity may be an important factor. While Hypersim contains more than 70K images, the images are collected from only 461 scenes. Many of them are variations of camera views and distances. In contrast, the synthetic images, while all of them are bedrooms, are all distinct scenes, which present diverse compositions of objects.
Blending Inputs and Outputs.
IADB [23] proposes a deterministic framework where the diffusion process is formulated as a series of interpolations between observations and noise. Although their training strategy produces deterministic mapping of observations and noise, the correlation between observation and noise in each pair is stochastic due to unpaired sampling during training. We analyze the applicability of this framework to deterministic dense prediction problems by sampling paired inputs and outputs and fine-tuning from a pre-trained T2I diffusion model. With such adaptation, the differences between their framework and our approach are only the variance schedule and parametrization, where the importance weight of inputs linearly rises through the diffusion process, and the U-Net predicts $y-x$.
Table 9 shows the comparison between DMP and IADB on surface normal estimation, and Table 10 is the result of depth estimation. Figure 11 demonstrates that the images generated by the model fine-tuned through IADB have noise and inaccurate predictions.
In addition to $\alpha$-blending used by DMP and IADB, we investigate the effect of an advanced blending strategy, Poission blending [51], which blends source and target images by solving a least-square fitting while reserving the gradient of source images. We assume image gradients are meaningful in the latent space. The diffusion process is viewed as increasing the intensity of the mask for selection editing. We adopt an off-the-shelf PyTorch implementation [4]. The performance on surface normal estimation is shown in Table 9, and the example outputs in Figure 11 show that the image quality is unsatisfying.
Table 9: Comparions with IADB [23] and Poission blending [51] on surface normal estimation.
In-domain Out-of-domain
[…]
IADB [23] 0.0675 0.1416 0.0974 0.2017
Poission [51] 0.0868 0.1888 0.1201 0.2623
DMP 0.0514 0.1156 0.0872 0.1886
Table 10: Comparions with IADB [23] on depth estimation.
In-domain Out-of-domain
[…]
IADB [23] 0.3099 0.4982 0.1165 0.5049 0.3132 0.1467
DMP 0.1072 0.8861 0.1020 0.2117 0.6395 0.1360
Figure 11: Qualitative comparisons between different blending frameworks.
B.4 Additional Comparison
Table 11: Quantitative comparisons with ControlNet [86] and Palette [60] on 3D property estimation and intrinsic image decomposition.
Normal Depth Albedo Shading
In Out In Out In Out In Out
[…]
ControlNet [86] 0.1021 0.2216 0.1862 0.4032 0.1739 0.7681 0.1287 0.4398 0.4004 0.2253 0.0302 0.0402 0.0265 0.0336
Palette [60] 0.1643 0.3642 0.1881 0.4160 0.6889 0.2626 0.3604 1.0535 0.2270 0.4203 0.0203 0.0199 0.0304 0.0260
DMP 0.0514 0.1156 0.0872 0.1886 0.1072 0.8861 0.0041 0.2117 0.6395 0.1360 0.0051 0.0064 0.1020 0.0070
Table 12: Quantitative comparisons with ControlNet [86] and Palette [60] on semantic segmentation.
Bed Pillow Lamp Window Painting Mean
[…]
ControlNet [86] 0.5215 0.4820 0.3540 0.1436 0.4275 0.2936 0.4999 0.4190 0.3823 0.3257 0.4370 0.3328
Palette [60] 0.0347 0.0329 0.0019 0.0018 0.0013 0.0012 0.0119 0.0119 0.0005 0.0005 0.0101 0.0097
DMP 0.8947 0.8506 0.5871 0.3645 0.6399 0.4414 0.8338 0.7335 0.7490 0.6735 0.7409 0.6127
ControlNet.
ControlNet [86] proposes a conditional text-to-image framework with additional control, such as edges or human poses, which constrains structures and layouts of output images. Since it is also an image-to-image generative model, we train it to take input images as control and output estimations. The performance of estimating 3D properties and intrinsic images is presented in Table 11, and the segmentation results are shown in Table 12. It demonstrates weaker generalizability than our approach.
In addition, we analyze the influence of varying initial noise in Figure 12. While the rough structures of the images are controlled by the input images, the initial noise alters the details of estimations. This variation is not tolerated for dense prediction.
Figure 12: Results of ControlNet with different initial noise. The outputs are not deterministic.
Palette.
Besides training GAN-based generative models from scratch and fine-tuning pre-trained diffusion models with the approaches listed in Section 4.1, we additionally include training an image-to-image diffusion model from scratch for comparison. Following the design of Palette [60], we expand the input layers of the U-Net to encode the concatenation of input and output images, with the U-Net parameterized to predict noise. The same autoencoder in the pre-trained diffusion model is also adopted. The performance is shown in Table 11 and Table 12, which indicates the inability of this approach to handle categorical label maps.
B.5 Improving Compared Methods
Since the results of GAN-based generative models consistently outperform diffusion-based models in our experiments, we seek performance enhancement for diffusion-based approaches. All experiments are conducted on in-domain surface normal estimation.
Figure 13: Qualitative comparisons of SDEdit with starting from different time steps. A trade-off exists between the effect of style transfer and content preservation.
Table 13: Quantitative comparisons on in-domain surface normal estimation between different timesteps where the generation process of SDEdit starts. The performance improves at the expense of deviation from input image contents.
[…]
SDEdit.
[…]
Figure 14: Qualitative comparisons between DDIB and DDIB with Plug-and-Play (PnP). The image contents are reserved but not consistent with accurate normals.
Table 14: Quantitative comparisons on in-domain surface normal estimation between DDIB and DDIB with Plug-and-Play (PnP). The feature injection regulates the generated contents while improving performance.
[…]
DDIB 0.1849 0.4210
DDIB + PnP 0.1652 0.3634
DDIB.
As presented in Section B.7, DDIB is capable of generating images that are likely sampled from output distributions, but the contents and geometry of output results are not consistent with input images. We explore the approach to content consistency by adopting feature constraints proposed by Plug-and-Play (PnP) [76] for image-to-image translation, which injects the self-attention and convolution features of input images into output images. As shown in Figure 14 and Table 14, the structures and contents of output images of DDIB with PnP constraints highly resemble the input images, but the estimated normals remain inaccurate despite better quantitative performance.
IP2P.
We analyze the expressiveness of inverted tokens by varying the number of training tokens in IP2P (learned). While the performance is slightly improved in one metric of quantitative evaluation, shown in Table 15, Figure 15 reveals that the differences between the estimated results are not significant.
Table 15: Quantitative comparisons on in-domain surface normal estimation between different training tokens of IP2P (learned). The increased number of tokens does not guarantee improved performance.
[…]
1 0.3550 0.7181
2 0.3470 0.7790
4 0.3274 0.6384
Figure 15: Qualitative comparisons of IP2P with different training tokens.
B.6 Failure Cases
We demonstrate some examples of failure cases in Figure 16 for surface normal estimation, Figure 17 for depth estimation, and Figure 18 for semantic segmentation, where off-the-shelf approaches might provide more accurate prediction than our method.
Table 16: Style templates, where {} is replaced by original prompts.
  • anime artwork, {} . anime style, key visual, vibrant, studio anime, highly detailed
  • concept art, {} . digital artwork, illustrative, painterly, matte painting, highly detailed
  • comic, {} . graphic illustration, comic art, graphic novel art, vibrant, highly detailed
  • neonpunk style, {} . cyberpunk, vaporwave, neon, vibes, vibrant, stunningly beautiful, crisp, detailed, sleek, ultramodern, magenta highlights, dark purple shadows, high contrast, cinematic, ultra-detailed, intricate, professional
    surrealist art, {} . dreamlike, mysterious, provocative, symbolic, intricate, detailed
  • abstract style, {} . non-representational, colors and shapes, expression of feelings, imaginative, highly detailed
  • art deco style, {} . geometric shapes, bold colors, luxurious, elegant, decorative, symmetrical, ornate, detailed
  • vaporwave style, {} . retro aesthetic, cyberpunk, vibrant, neon colors, vintage 80s and 90s style, highly detailed
Figure 16: Failure cases of surface normal estimation.
Figure 17: Failure cases of depth estimation.
Figure 18: Failure cases of semantic segmentation.
B.7 Results of Compared Methods
We show the example images generated by the compared methods listed in Section 4.1. The results of surface normal estimation are in Figure 19, with depths in Figure 20, albedo in Figure 21, shading in Figure 22, and semantic segmentation in Figure 23.
Figure 19: Qualitative results of compared methods on surface normal estimation.
Figure 20: Qualitative results of compared methods on depth estimation.
Figure 21: Qualitative results of compared methods on albedo estimation.
Figure 22: Qualitative results of compared methods on shading estimation.
Figure 23: Qualitative results of compared methods on semantic segmentation.
Appendix C Implementation Details
Model Architecture and Optimization.
We use Stable Diffusion 1.4 as the pre-trained text-to-image model and adapt it with rank $=4$ for LoRA. We fine-tune the model for $50$K steps with batch size $8$ and learning rate $0.0001$ with a cosine decay schedule. The training takes around 14 hours with a single NVIDIA RTX 3090.
Generating Images.
We generate the training and test images by first generating a set of prompts with a large language model. The prompt for the language model is a template adapted from pix2pix-zero [50], where different scene keywords are filled in. The template is
“Provide a caption for a photo of a scene. The caption should contain many adjectives, should describe colors, styles, lighting and materials in the photo, should be in English and should be no longer than 150 characters. Caption:”.
The placeholder scene is replaced by “bedroom” for training images and in-domain test images. To generate out-of-domain test images for estimating 3D properties and intrinsic images, it is replaced by uniform sampling from the keywords in LABEL:tab:scenes.
Out-of-domain test images for segmentation are synthesized by varying the image styles of in-domain test images, for semantic categories should remain the same across training and test images. The prompts regulating the styles are listed in Table 16 borrowed from an online post [1].
Appendix D Applications
Surface normals and depths facilitate many vision tasks. We show by the examples of 3D photo inpainting [65] that precise depths improve 3D reconstruction from 2D images. Compared to the default depth estimator [54], the resulting videos produced with the depth maps generated by our approach have more accurate depth relationships between the objects. Please refer to the project website for visual demonstrations.
Table 17: Scenes categories of out-of-domain images.
airlock airplane cabin airport terminal airport ticket counter
alcove amusement arcade anechoic chamber indoor apse
aquarium arcade archive armory
indoor arrival gate art gallery art school art studio
artists loft assembly line indoor athletic field attic
auditorium auto factory indoor auto mechanics auto showroom
backstage indoor badminton court baggage claim ball pit
ballroom indoor bank bank vault banquet hall
indoor baptistry bar barbershop barrack
basement indoor basketball court bathhouse bathroom
indoor batting cage indoor bazaar beauty salon bedchamber
bedroom beer hall belfry bell foundry
berth berth deck betting shop bicycle racks
bindery biology laboratory indoor bistro indoor bleachers
indoor bomb shelter bookbindery bookstore indoor booth
indoor bow window bowling alley box seat boxing ring
breakroom indoor brewery indoor brickyard burial chamber
indoor bus depot bus interior indoor bus station butchers shop
indoor cabin cafeteria call center candy store
canteen backseat car interior frontseat car interior cardroom
cargo container interior indoor carport indoor casino catacomb
indoor cathedral catwalk chapel checkout counter
cheese factory chemistry lab indoor chicken coop indoor chicken farm
childs room interior choir loft indoor church indoor circus tent
classroom clean room indoor clock tower indoor cloister
closet clothing store cockpit coffee shop
computer room conference center conference hall conference room
confessional control room indoor control tower indoor convenience store
corridor courtroom interior covered bridge crawl space
cybercafe indoor dairy dance school darkroom
day care center delicatessen dentists office department store
departure lounge indoor diner dining car dining hall
dining room discotheque distillery indoor doorway
dorm room dress shop dressing room indoor driving range
drugstore editing room electrical room elevated catwalk
interior elevator elevator lobby elevator shaft engine room
entrance hall indoor escalator exhibition hall fabric store
indoor factory fastfood restaurant indoor ferryboat indoor firing range
fishmarket interior fitting room indoor flea market indoor florist shop
food court indoor foundry funeral chapel funeral home
furnace room galley game room indoor garage
indoor general store indoor geodesic dome gift shop great hall
indoor greenhouse indoor gun deck gun store indoor gymnasium
hallway indoor hangar hardware store hat shop
hatchery hatchway hayloft hearth
home office home theater hospital room indoor hot tub
hotel breakfast area hotel room indoor hunting lodge ice cream parlor
indoor ice skating rink indoor incinerator indoor inn indoor jacuzzi
indoor jail jail cell jewelry shop jury box
indoor kennel kindergarden classroom indoor kiosk kitchen
kitchenette lab classroom indoor labyrinth landing
laundromat lavatory lecture room legislative chamber
indoor library indoor lido deck limousine interior indoor liquor store
living room lobby locker room loft
indoor lookout station indoor lumberyard machine shop indoor market
martial arts gym maternity ward mess hall mezzanine
military hospital mill mine indoor mini golf course
indoor monastery morgue indoor mosque indoor movie theater
indoor museum music store music studio natural history museum
newsroom indoor newsstand nightclub indoor nuclear power plant
nursery nursing home indoor observatory office
office cubicles indoor oil refinery operating room optician
orchestra pit interior organ loft orlop deck ossuary
indoor outhouse oyster bar packaging plant palace hall
pantry paper mill indoor parking garage parlor
particle accelerator indoor party tent pawnshop penalty box
perfume shop pet shop pharmacy physics laboratory
piano store pig farm indoor pilothouse pizzeria
indoor planetarium playroom indoor podium portrait studio
indoor power plant print shop promenade deck indoor pub
pulpit pump room indoor quonset hut reading room
reception recreation room indoor recycling plant refectory
repair shop restaurant restaurant kitchen indoor restroom
revolving door riding arena indoor roller skating rink rolling mill
sacristy sauna sawmill science museum
scriptorium security check point server room sewer
sewing room shipping room indoor shipyard shoe shop
indoor shopping mall shower shower room shrine
indoor skywalk sporting goods store squash court stable
indoor stage staircase indoor steam plant indoor steel mill
storage room storeroom food submarine interior subway interior
supermarket sushi bar indoor swimming pool indoor synagogue
tearoom teashop television room television studio
indoor tennis court indoor tent textile mill indoor procenium theater
indoor round theater indoor seats theater thriftshop throne room
ticket booth indoor ticket window indoor tobacco shop toyshop
indoor track trading floor train interior rail indoor tunnel
road indoor tunnel turkish bath utility room utility tunnel
van interior indoor velodrome ventilation shaft vestry
veterinarians office videostore indoor volleyball court voting booth
waiting room walk in freezer indoor warehouse indoor washhouse
indoor water treatment plant wet bar whispering gallery wig shop
window seat winery witness stand workroom
workshop indoor wrestling ring youth hostel basketball arena
football arena hockey arena performance arena rodeo arena
soccer arena home atrium public atrium bakery kitchen
bakery shop airplane cargo deck choir loft cloakroom booth
cloakroom library cubicle office cubicle home dinette
vehicle dinette elevator door freight elevator ferryboat cargo deck
fitting room organ loft establishment poolroom home poolroom
spa massage room spa mineral bath corridor in a subway station platform in a subway station
turnstiles in a subway station platform in a train station station in a train station barrel storage in a wine cellar
bottle storage in a wine cellar